find the equation of one of the curves whose slope at any point is equal to y + 2x .

Grade:

1 Answers

SHAIK AASIF AHAMED
9 years ago
given y1=y+2x
y1-y=2x
I.F=e-x
ye-x=\int 2xe^{-x}
on solving we get
y=-2x-2+cex
Thanks and Regards
Shaik Aasif
askIITians Faculty